solve system of linear equation by elimination
a+9b=1
a+3b=-5

Answer by mananth(16946)

a + 9 b = 1 .............1

a + 3 b = -5 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -3
1 a 9 b = 1
-3 a -9 b = 15
Add the two equations
-2 a = 16
/ -2
a = -8
plug value of a in (1)
1 a + 9 b = 1
-8 + 9 b = 1
9 b = 1 + 8
9 b = 9
b = 1
a -8
b 1
m.ananth@hotmail.ca